# Understanding Figurative Meaning through Explainable Visual Entailment

### v2

@@ F.4.1 Few-shot Prompt for generating entailing claims
F.4.2 Zero-shot Prompt for validating the entailing captions
@@
F.4.3 Few-shot Prompt for generating opposite claims
@@
F.4.4 Zero-shot Prompt for generating explanations
@@
Appendix G Model Taxonomy
@@
Appendix H Multimodal Structured Chain-of-Thought Performance
@@
Appendix I Additional Models
@@
Appendix J By-Phenomenon Performance
@@
Appendix K Human Baseline
| Phenomenon | Dataset |
| Metaphor /Similes | HAIVMET | 78.84 | 81.25 | ||||
94.36 | 77.78 | |||||
| Idioms |
89.26 | 49.74 | ||||
| Sarcasm | MuSE | 68.89 | 85.42 | ||||
| Humor | MemeCap | 100.0 | 78.03 | ||||
| NYCartoons | 71.43 | 47.83 | |||||
| Overall | 89.09 | 77.26 | |||||
Appendix L How Well Does the Explanation Score Predict Human Judgment on Adequacy?
We explore whether the proposed explanation score can capture human judgement of explanation adequacy. We collect all instances from Section 5 where both annotators agreed on the adequacy judgement for the explanation. We evaluate if the explanation score described in Section 4.2 can act as a good predictor for the human adequacy judgment. We find that the area under the Precision-Recall curve is 0.79, and the maximum F1 score is 0.77, obtainable at the explanation score threshold of 0.53. Hence, we use this threshold to report the results in Table 3. We also use the threshold of 0.6 since it maximizes F1 such that both precision and recall are above 0.75.
Appendix M How Do Models Perform When Only Predicting the Label?
@@
Appendix N Annotation Interfaces
@@
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/vismet_Ex2.png)
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/irfl_example_resized.jpeg)
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/muse_ex_resized.jpeg)
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/memecap_ex_resized.jpeg)
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/ny_ex_resized.jpeg)

### v3

@@ E.4.1 Few-shot Prompt for generating entailing claims
E.4.2 Zero-shot Prompt for validating the entailing captions
@@
E.4.3 Few-shot Prompt for generating opposite claims
@@
E.4.4 Zero-shot Prompt for generating explanations
@@
Appendix F Model Taxonomy
@@
Appendix G Multimodal Structured Chain-of-Thought Performance
@@
Appendix H Additional Models
@@
Appendix I By-Phenomenon Performance
@@
Appendix J How Do Models Perform When Only Predicting the Label?
@@
Appendix K Annotation Interfaces
@@
![[Uncaptioned image]](extracted/6211238/figs/dataset_exs/vismet_Ex2.png)
![[Uncaptioned image]](extracted/6211238/figs/dataset_exs/irfl_example_resized.jpeg)
![[Uncaptioned image]](extracted/6211238/figs/dataset_exs/muse_ex_resized.jpeg)
![[Uncaptioned image]](extracted/6211238/figs/dataset_exs/memecap_ex_resized.jpeg)
![[Uncaptioned image]](extracted/6211238/figs/dataset_exs/ny_ex_resized.jpeg)